OpenStudy (anonymous):

2 log (base 2) (8x) evaluate

OpenStudy (ybarrap):

2 log (8x), base 2 2 * (log 8 + log x) = 2 * (3 + log x), log 8 = 3 since 2^3 = 8 = 6 + 2*logx